Wild type 710

green 70

dwarf 140

dwarf, curly leaves 90

green, dwarf, curly leaves 680

green, curly leaves 150

Determine order and distance between genes
Answer #1

Answer:

Wild type 710 (+ + +)

green 70 (g + +)

dwarf 140 (+ d +)

dwarf, curly leaves 90 (+ d c)

green, dwarf, curly leaves 680 (g d c)

green, curly leaves 150 (g + c)

Total = 1840

Hint: always parental combinations are more numbered than any other recombinants in linkage. Hence, parental trihybrid would be +++/gdc.

1). If single cross over occurs between g&d...

Normal : ++/gd

After cross over: + d/g +

+ d progeny= 140+90=230

g + progeny= 150+70=220

Total of this progeny= 450

Recombination frequency between g & d = (no. of recombinants/total)100

= (450/1840)100= 24.46%

2). If single cross over occurs between d and c...

Normal: ++/dc

After cross over: +c/d+

+c progeny= 150

d + progeny= 140

Total of this progeny= 290

Recombination frequency between d&c= (290/1840)100

= 15.76%

3). If single cross over occurs between g and c...

Normal: ++/gc

After cross over: +c/g+

+c progeny= 90

g + progeny= 70

Total of this progeny= 160

Recombination frequency between g&c= (160/1840)100= 8.7%

Recombination frequency (%) = Distance between genes (m.u.)

Order of genes = g------8.7m.u.-----c----15.76m.u.-------d

Gene order= g----c---d